How long would it take to triple an investment at simple interest with a rate of 4.5% per year
Answer by stanbon(75887)

How long would it take to triple an investment at simple interest with a rate of 4.5% per year
------
3P = P(1+rt)
------
1+rt = 3
------
rt = 2
0.045t = 2
time = 2/(0.045) = 44.44 years
-------------------
Cheers,
Stan H.
-------------